OpenStudy (summer4ever):

2x + 4y = 12 3x - 5y = 10 Question: A. Write the equations in slope-intercept form. (Show your work.) B. Graph the pair of linear equations. C. Use the graph to estimate the solution to the system of equations.

OpenStudy (summer4ever):

@Mythorius

alones (alones):

tag green users there smart or tag @lυἶცἶ0210 lel

OpenStudy (summer4ever):

@lυἶცἶ0210

alones (alones):

no dude don't tag him..tag green users there more active

OpenStudy (haleyelizabeth2017):

Slope intercept form is y=mx+b, can you try to turn: 2x + 4y = 12 3x - 5y = 10 into slop intercept form? step 1: subtract x from both sides.

OpenStudy (summer4ever):

then what? i dont really understand it. im doing a benchmark for k12 and nly have 5 more minutes to submit it or else i get a 0

OpenStudy (haleyelizabeth2017):

2x + 4y = 12 3x - 5y = 10 I will do the first one for you, and then you need to do the second one. 2x+4y=12 subtract x from both sides -2x -2x 4y=-2x+12 divide both sides by the integer that is multiplied by y /4 /4 y=-2/4x+3 y is now isolated y=-1/2x+3 simplify

OpenStudy (haleyelizabeth2017):

Now you try the second one, one step at a time (: I will make sure it is correct

OpenStudy (summer4ever):

-x+2y=6?

OpenStudy (haleyelizabeth2017):

No...read through my steps again. Subtract the x from both sides

OpenStudy (haleyelizabeth2017):

Well, subtract 3x from both sides...

OpenStudy (summer4ever):

so the one you just did for me when you said symplify is that first one completely solved so i can graph it?

OpenStudy (haleyelizabeth2017):

Yes.

OpenStudy (summer4ever):

this is the graph

OpenStudy (haleyelizabeth2017):

But you need to get the second one done, so we can graph both

OpenStudy (summer4ever):

3x - 5y = 10 -3 -3

OpenStudy (summer4ever):

good so far?

OpenStudy (haleyelizabeth2017):

-3x

OpenStudy (haleyelizabeth2017):

But yes

OpenStudy (summer4ever):

ok so -3x - 5y = 10-3x=

OpenStudy (haleyelizabeth2017):

I mean divide both sides by -5 lol

OpenStudy (haleyelizabeth2017):

http://prntscr.com/accx4f

OpenStudy (summer4ever):

so -3x/-5?

OpenStudy (haleyelizabeth2017):

Yes, which is 3/5x so that one is y=3/5x-2 like Gaby said

OpenStudy (summer4ever):

okay thank you

OpenStudy (haleyelizabeth2017):

No problem.